Kunle tie his mother's wrapper round his waist completely 12times. *If the measurement of his waist has a diameter 3.5cm. Find the length of the wrapper*
1
Expert's answer
2020-06-08T19:05:52-0400

Let "L=" the length of the wrapper, "d=" the diameter of his waist, and "n="number of wraps.

Given "d=3.5\ cm, n=12."

Then


"L=\pi\cdot d\cdot n=\pi\cdot3.5\ cm\cdot12=42\pi\ cm\approx132\ cm"

The length of the wrapper 132 cm.